# ScholaWrite: A Dataset of End-to-End Scholarly Writing Process

### v2

@@ 4 ScholaWrite: A Dataset of Cognitive Process of Scholarly Writing
| 1 | 2 | 3 | 4 | 5 | |
| Idea Generation | 515 | 130 | 116 | 309 | 3255 |
| Idea Organization | 0 | 45 | 25 | 9 | 231 |
| Section Planning | 182 | 57 | 111 | 201 | 773 |
| Text Production | 9267 | 2438 | 5109 | 4478 | 14031 |
| Object Insertion | 583 | 383 | 62 | 486 | 1300 |
| Cross-reference | 141 | 112 | 13 | 292 | 458 |
| Citation Integration | 75 | 151 | 69 | 127 | 245 |
| Macro Insertion | 16 | 7 | 51 | 29 | 33 |
| Linguistic Style | 233 | 75 | 42 | 201 | 411 |
| Coherence | 422 | 242 | 126 | 193 | 1021 |
| Clarity | 1249 | 645 | 721 | 1180 | 3301 |
| Scientific Accuracy | 307 | 15 | 2 | 24 | 95 |
| Structural | 359 | 506 | 105 | 257 | 1042 |
| Fluency | 116 | 90 | 46 | 135 | 476 |
| Visual Formatting | 752 | 163 | 43 | 427 | 567 |
Summary
Our post-processed dataset, ScholaWrite, contains end-to-end writing trajectories with intentions annotated, for five Overleaf projects whose final product turned into arXiv preprints. The dataset consists of 61,504 arrays of keystroke changes with fine-grained annotations of writing intentions. We publicly release our post-processed data in Huggingface data cards.888https://huggingface.co/datasets/minnesotanlp/scholawrite Table 3 shows the overall statistics of the dataset.
Writing Intention Distributions
The writing intention distributions is illustrated in Figure 5(a). During the planning stage, idea generation is the most frequent intention. Text production dominates the implementation stage, and clarity becomes the primary focus during the revision stage. Since we defined the boundary between implementation and revision as the point where meaning changes occur, the high frequency of text production suggests that authors often continue revising their texts by introducing new topics into the documents. Table 3 shows the label distribution per project.
Flows Between Writing Intentions
We analyze the flow of writing intentions most likely to follow each other throughout the writing process, as presented in Table 4. Many intentions proceed to text production, where the idea generation has the highest probability of feeding into text production. Also, the reflexive relationship between text production and clarity suggests a constant loop of producing new texts and refining them, ensuring the precision and reasonability of their claims.
@@
Time-series Distributions of Writing Intentions
We calculated the average Wasserstein distance between each intention distribution and uniform distribution to assess how evenly intentions occur throughout the writing process. As shown in Figure 5(b), text production had the smallest distance, while macro insertion had the largest. This suggests that text production is spread throughout the writing, whereas macro insertion is concentrated in specific phases. A similar pattern appears in Figure 5(c), where text production is evenly distributed over time, and macro insertion peaks in the early and final stages of the writing process. Figures 6(a) and 6(b) demonstrate a similar pattern that text production of the implementation stage spreads evenly over time. Types of clarity and structural revisions tend to occur after the medium steps of writing. Please refer to Figures 9 to 12 in the Appendix for each of the five Overleaf projects.
5 Applications for Writing Assistance with ScholaWrite
We believe ScholaWrite can serve as a valuable resource for training language models or enhancing the cognitive writing process in future writing assistants for scholarly writing. To evaluate its usability, we performed several experiments where existing LLMs were trained to mimic the complex, non-linear scientific writing processes of human scholars. Specifically, we aimed to showcase the capabilities of LLMs trained on ScholaWrite in two scenarios:
(1) Predicting an author’s next writing intention (Section 5.1)
: The task is crucial for writing assistants to accurately assess the writer’s current status in context and predict the correct writing intention. This enables them to offer cognitively-appropriate writing suggestions that align with the writer’s needs.
(2) Iteratively generating scholarly writing actions from scratch (called Iterative Self-Writing), mirroring the human writing process (Section 5.2)
: This task focuses on how well the model trained on our dataset can replicate the actual iterative writing and thinking process of scholars, and whether the generated text achieves higher quality compared to LLM-prompted writing.
As illustrated in Figure 7, the first task (represented by the inner box labeled "Prediction") takes the "before" text from a keystroke pair and context prompt, and predicts the writing intention to apply for the subsequent actions. Once the intention is predicted, the second task generates the “after” text based on the predicted intention and given prompt (represented by the outer box labeled “Generation”). This process repeats iteratively until no further changes are made or a maximum number of iterations, such as 100, is reached.
We randomly split our dataset into train (80%) and test (20%) sets. For each intention label in the test set, we randomly select up to 300 keystroke entries, due to budget constraints. Note that the same train and test sets are applied to all models across all experiments. See Appendix D for a full description of the model training process.
5.1 Predicting Next Writing Intention
Setup & Metrics
We use BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and Llama3.1-8B-Instruct Dubey et al. (2024) as baselines, fine-tuning each on the ScholaWrite training set. For comparison, we also run GPT-4o OpenAI (2024) on the test set. The fine-tuning prompt included all possible labels with definitions, task instructions, the “before” text chunk, and the corresponding human-annotated intention label, asking the model to predict the intention label based on the “before-text”. Differences in prompts were limited to only task instructions (see Appendix D.3 for prompt details). To evaluate model performance, we used a weighted f-1 score999Weighted F-1 was chosen to address skewed label distribution (as shown in Figure 5(a) and Table 1)., comparing predictions to gold intention labels on the test set.
| BERT | RoBERTa | LLama-8B | GPT-4o | |
| Base | 0.04 | 0.02 | 0.12 | 0.08 |
| + SW | 0.64 | 0.64 | 0.13 | - |
Results
Table 5 presents the weighted F1 scores for predicting writing intentions across baselines and fine-tuned models. Regardless of the intricate nature of the task itself101010Each model is asked to predict the next intention by only giving the ‘before’ text. Human annotators, however, label the data by looking through multiple consecutive ‘before and after’ edits to determine the current intention. In addition, the next intention chosen by the author does not necessarily mean that it is the only correct intention., all models finetuned on ScholaWrite show an improved performance compared to their baselines. BERT and RoBERTa achieved the most improvement, while LLama-8B-Instruct showed a modest improvement after fine-tuning. As detailed in Appendix D, more training epochs and encoder-decoder architectures of BERT variants are assumed to be the reason for significant improvement compared to LLMs. This aligns with findings from Grasso and Locci (2024); Yu et al. (2023), which shows that RoBERTa and BERT can often match or even outperform LLMs for the text classification tasks. Those results demonstrate the effectiveness of our ScholaWrite dataset to align language models with writers’ intentions.
5.2 Iterative Self-Writing
Setups
Iterative self-writing involves two subtasks (Figure 7): (1) next intention prediction and (2) "after-text" generation. For intention prediction, we use fine-tuned Llama3.1-8B-Instruct on the ScholaWrite training set (“Llama-8B-SW-pred”), compared to vanilla models (Llama-8B-Instruct, and GPT-4o) from Section 5.1 as baselines. Please refer to Figure 13 in Appendix D for the detailed pipeline.
For “after-text” generation, another vanilla Llama3.1-8B-Instruct is fine-tuned on the ScholaWrite training set (“Llama-8B-SW-gen”), with baselines including vanilla Llama3.1-8B-Instruct and GPT-4o. The fine-tuning prompt included task instructions, a verbalizer derived from human-annotated labels, and the “before” text. The model’s task is to generate “after-text” given the verbalizer and “before” text. To maintain consistency across outputs, prompts were adjusted for each model, with differences mainly in task instructions (see Appendix D.3 for the prompt templates).
Seed documents were derived from four award-winning NLP papers spanning different topics Zeng et al. (2024); Lu et al. (2024b); Du et al. (2022a); Etxaniz et al. (2024). Listing LABEL:table:seed-entry presents the LaTeX-formatted abstract of Du et al. (2022a) as an example seed text for initiating the iterative self-writing experiment.
6 Conclusion and Future Work
We introduce ScholaWrite, the first-of-its-kind dataset that captures the end-to-end scholarly writing process, consisting of nearly 62K LaTeX keystrokes collected via our custom-built Chrome extension. This dataset, sourced from ten graduate students with varying levels of scientific writing expertise, is further enriched with expert annotations based on a novel taxonomy of cognitive writing intentions inspired by Flower and Hayes (1981); Koo et al. (2023). Our study demonstrates the value of ScholaWrite as a resource for advancing the cognitive capabilities of large language models (LLMs) and developing cognitively-aligned writing assistants, enabling them to emulate the complex, non-linear, and intention-driven nature of human writing.
Limitations and Ethical Considerations
Acknowledgements
This work was supported by the research gift from Grammarly. We thank Anna Martin-Boyle and Ryan Koo for their insights and the data collection tool based on their earlier version of this paper. We thank the participants of our data collection for their valuable time, effort, and contributions, which were essential to the success of this research. We also thank Minnesota NLP group members and anonymous reviewers for providing us with valuable feedback and comments on the paper draft.
Appendix A More About Data Collection Process
A.1 Participant Recruitment & Demographics
A.2 Technical Details of System Implementations
A.3 Data Post-Processing
[…]
Appendix B More About Writing Intention Annotations & Taxonomy
B.1 Annotator Recruitment
B.2 Detailed Annotation Process
Appendix C More About the Taxonomy of Scholarly Writing Process
During the planning stage, the writers engage in a process of generating and organizing raw ideas, arguments, or content structures that were not introduced in the previous trajectory. Based on the plan, the writers implement their plan by drafting full sentences and paragraphs and structuring the contents tangibly. At the same time, the writers enter the revision stage by improving the quality of their implemented sentences and LaTeX objects in terms of linguistic styles, format, or information accuracy. Particularly, spans of keystrokes whose intentions involved any changes but did not change the meaning of original texts are classified as Revision. For those edits that show changes in the meaning, we considered them as Implementation. Furthermore, if an author repeatedly adds, removes, and revises text back and forth until a sentence is completed, we consider this process as part of text production. Any subsequent changes made to the sentence after it is finished are considered revision.
Appendix D More about ScholaWrite Evaluation
D.1 Training environments
@@
D.2 Details About Finetuning Process
@@
The fine-tuning prompt consisted of task instructions, a list of all intention labels with their definitions, the “before” text, and the corresponding human-annotated intention label. To achieve optimal performance while minimizing memory usage, we employed QLoRA Dettmers et al. (2024) to fine-tune all linear modules of a 4-bit quantized Llama. During fine-tuning, we utilized the ‘train_on_response_only’ function provided by the unsloth library, which masks the task instructions, intention label definitions, and “before” text with -100s. This ensures the model is trained exclusively on the response portion of the fine-tuning prompt (i.e., the predicted intention label), without being influenced by the instructional components of the input. The model was fine-tuned for one epoch with a batch size of 2, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
@@
D.3 Fine-tuning Prompt Templates
D.3.1 Prediction Prompt for Llama-8B-Zero models
@@
D.3.2 Generation Prompt for Llama-8B-Zero models
@@
D.3.3 Prediction Prompt for Llama-8B-SW models
@@
D.3.4 Generation Prompt for Llama-8B-SW models
@@
## Input: before_text
## Output:
D.3.5 Classification Prompt for GPT-4o
“You are a classifier that identify the most likely next writing intention. You will be given a list of all possible writing intention labels with definitions, and an in-progress LaTex paper draft written by a graduate student. Please strictly follow user’s instruction to identify the most likely next writing intention
Here is a list of all the possible writing intention labels with definitions:
@@ D.3.6 Generation Prompt for GPT-4o
D.4 Seed Documents for Iterative Self-Writing

### v3

@@ Appendix D ScholaWrite dataset statistics
1 | 2 | 3 | 4 | 5 | |
Idea Generation | 515 | 130 | 116 | 309 | 3255 |
Idea Organization | 0 | 45 | 25 | 9 | 231 |
Section Planning | 182 | 57 | 111 | 201 | 773 |
Text Production | 9267 | 2438 | 5109 | 4478 | 14031 |
Object Insertion | 583 | 383 | 62 | 486 | 1300 |
Cross-reference | 141 | 112 | 13 | 292 | 458 |
Citation Integration | 75 | 151 | 69 | 127 | 245 |
Macro Insertion | 16 | 7 | 51 | 29 | 33 |
Linguistic Style | 233 | 75 | 42 | 201 | 411 |
Coherence | 422 | 242 | 126 | 193 | 1021 |
Clarity | 1249 | 645 | 721 | 1180 | 3301 |
Scientific Accuracy | 307 | 15 | 2 | 24 | 95 |
Structural | 359 | 506 | 105 | 257 | 1042 |
Fluency | 116 | 90 | 46 | 135 | 476 |
Visual Formatting | 752 | 163 | 43 | 427 | 567 |
Figures 9 to 12 show several characteristics of human writing process, analyzed from ScholaWrite dataset: (1) Figure 9 - the average Wasserstein distance between each intention distribution and uniform distribution; (2) Figure 10 - distribution of labels over time; (3) Figure 11 for high-level intention distribution over time; and (4) Figure 12 for intention-wise writing activity distribution over time.
Appendix E More about ScholaWrite Evaluation
E.1 Training environments
@@
E.2 Details About Finetuning Process
@@
The fine-tuning prompt included all possible labels with definitions, task instructions, the “before-text” chunk, and the corresponding human-annotated intention label, asking the model to predict the intention label based on the “before-text”. Differences in prompts were limited to only task instructions (see §E.3 for prompt details).
To achieve optimal performance while minimizing memory usage, we employed QLoRA Dettmers et al. (2024) to fine-tune all linear modules of a 4-bit quantized Llama. During fine-tuning, we utilized the ‘train_on_response_only’ function provided by the unsloth library, which masks the task instructions, intention label definitions, and “before” text with -100s. This ensures the model is trained exclusively on the response portion of the fine-tuning prompt (i.e., the predicted intention label), without being influenced by the instructional components of the input. The model was fine-tuned for one epoch with a batch size of 2, 4 gradient accumulation steps, and the AdamW 8-bit optimizer.
@@
For finetuning, we randomly split the ScholaWrite dataset into training (80%) and testing (20%) sets. From each intention label in the test set, we sample up to 300 keystroke entries due to budget constraints.
For intention prediction, we fine-tune Llama3.1-8B-Instruct on ScholaWrite training set (Llama-8B-SW-pred”) and compare it to baseline models (Llama-8B-Instruct and GPT-4o) from Section 5.1. For “after-text” generation, we fine-tune another Llama3.1-8B-Instruct model (“Llama-8B-SW-gen”) using the same dataset, with Llama3.1-8B-Instruct and GPT-4o as baselines. The fine-tuning prompt includes task instructions, a verbalizer from human-annotated labels, and “before-text.” While prompts were standardized, task instructions varied by model.
@@
Also, due to budget constraints, models had different revision strategies. Llama-8B-SW-* and Llama-8B-Instruct continued revision until the next predicted intention changed. GPT-4o, however, moved to the next iteration regardless. We refer to the fine-tuned Llama-8B model as Llama-8B-ScholaWrite (or Llama-8B-SW) and the vanilla model as Llama-8B-Zero.
E.3 Prompt Templates
E.3.1 Prediction Prompt for Llama-8B-Zero models
@@
E.3.2 Generation Prompt for Llama-8B-Zero models
@@
E.3.3 Prediction Prompt for Llama-8B-SW models
@@
E.3.4 Generation Prompt for Llama-8B-Zero models
@@
## Input: before_text ”
E.3.5 Classification Prompt for GPT-4o
“You are a classifier that identify the most likely next writing intention. You will be given a list of all possible writing intention labels with definitions, and an in-progress LaTex paper draft written by a graduate student. Please strictly follow user’s instruction to identify the most likely next writing intention.
Here are the verbalizers of all the possible writing intention labels:
@@ E.3.6 Generation Prompt for GPT-4o
E.4 Seed Documents for Iterative Self-Writing